Write a quadratic equation with -5 abnd 2/3 as its roots. Write the equation in the form ax^2+bx+c=0
(x-5)(x+2/3) is what I have gotten so far. I am not sure if I am heading in the right direction.

Answer by lwsshak3(11628)

Write a quadratic equation with -5 abnd 2/3 as its roots. Write the equation in the form ax^2+bx+c=0
***
(x+5)(3x-2)=3x^2+13x-10=0